OpenStudy (anonymous):

Using the graph, which of these ordered pairs is the best estimate of the solution of the system y = 3x + 2 and y = -1/4 x + 1? graphs of y = 3x minus 1, y=negative one-fourth x + 1, y=negative two-thirds x, y = 3x + 2 A. (0.5, 0.5) B. (-0.3, 1.1) C. (1.2, 0.8) D. (0.8, -0.5)

OpenStudy (anonymous):

Solve the system 3x - 5y = 8 and 4x - 7y = 10.

OpenStudy (anonymous):

Is the system consistent or inconsistent? y = -1/2x - 1 and y = 3x + 3